Titus Company produced 3,100 units of a product that required 3.779 standard hours per unit. The standard fixed overhead cost per unit is $2.60 per hour at 11,900 hours, which is 100% of normal capacity. Determine the fixed factory overhead volume variance. Enter a favorable variance as a negative number. Round your answer to the nearest dollar.
Answer #1
Standard hours for actual production 11715 =3100*3.779
Budgeted fixed overhead 30940 =11900*2.60
Less: Fixed overhead applied 30459 =11715*2.60
Fixed factory overhead volume variance 481 Unfavorable